Assume 300 students attend a dance. Which band would you choose and the cost per ticket you need to cover your expenses.
Band A charges $600 to play for the entire evening
Band B charges $350 plus $1.25 for each ticket sold

Found 2 solutions by stanbon, palanisamy:
Answer by stanbon(75887)

Assume 300 students attend a dance. Which band would you choose and the cost per ticket you need to cover your expenses.
Band A charges $600 to play for the entire evening
Flat Rate for any number of students: $600
with 300 students the charge will be $600
Cost per ticket to cover expenses would be 600/300 = $2.00
----------------------------------------------------------
Band B charges $350 plus $1.25 for each ticket sold
Band B charges $(350+1.25x) for "x" students
with 300 students the charge will be 350+1.25*300 = $725
Cost per ticket to cover expenses would be 725/300 = $2.43
----------------------------------------------------------
Cheers,
Stan H.

Answer by palanisamy(496)

300 students attend a dance.
Band A charges $600 to play for the entire evening.
Band B charges $350 plus $1.25 for each ticket sold
Charges for band B = 350+300*(1.25)
= 350+375
= 725 $
I will choose band A
and the the cost per ticket you need to cover the expenses = 600/300 = 2$